### v2

# Optimizing a Worldwide-Scale Shipper Transportation Planning in a Carmaker Inbound Supply Chain

Abstract
We study the shipper-side design of large-scale inbound transportation networks, motivated by Renault’s global supply chain. We introduce the Shipper Transportation Design Problem, which integrates consolidation, routing, and regularity constraints, and propose a tailored Iterated Local Search (ILS) metaheuristic. The algorithm combines large-neighborhood search with MILP-based perturbations and exploits bundle-specific decompositions and giant container bounds to obtain scalable lower bounds and effective benchmarks. Computational experiments on real industrial data show that the ILS achieves an average gap of 7.9% to the best available lower bound on world-scale instances with more than 700,000 commodities and 1,200,000 arcs, delivering solutions showing a potential of 23.2% cost reduction compared to the Renault-based benchmark. To our knowledge, this is the first approach to solve shipper-side transportation design problems at such scale. Our analysis further yields managerial insights: accurate bin-packing models are essential for realistic consolidation, highly regular plans offer the best balance between cost and operational stability, and outsourcing is only attractive in low-volume contexts, while large-scale networks benefit from in-house planning.
1 Introduction
Designing an efficient inbound supply chain is a critical task for manufacturing firms that operate at massive scale. The inbound supply chain connects a company’s suppliers to its production sites, and its configuration must often be revised in response to product introductions, demand shifts, carrier contracting, or the need to mitigate disruptions. In industries with globalized supplier bases and geographically distributed industrial sites, this task becomes especially complex. While many firms outsource logistics operations to third-party logistics providers (3PLs) to leverage shared transportation assets and reduce operational complexity, this outsourcing approach is not universally applicable. Some manufacturers operate supply chains of such magnitude and intricacy that they prefer to retain full control over their logistics planning. This is particularly true for large automotive manufacturers, who can achieve economies of scale internally. For these firms, outsourcing would not only reduce flexibility and transparency but also forgo substantial cost savings and introduce operational risk.
Renault exemplifies such a case. As a global car manufacturer with a vast supplier base and worldwide production footprint, Renault moves millions of parts across continents each year. This results in an inbound logistics network of exceptional size and complexity—comprising thousands of suppliers, hundreds of logistics platforms, and multiple industrial sites—requiring highly detailed and scalable optimization methods to ensure efficient operations. In this context, the company cannot rely on off-the-shelf tools or generic outsourcing models. Instead, it requires bespoke, high-performance optimization tailored to its specific needs.
This paper arises from a close industrial collaboration with Renault and addresses a real-world problem of strategic importance: the long-term planning of the company’s inbound transportation network. In this context, our objective is to produce a provisional transportation plan for parts to be assembled in the next six months. Each commodity—defined as a part to be delivered from a specific supplier to a specific industrial site by a given date—must be routed through Renault’s logistics network in a cost-efficient manner. The decision involves choosing whether to ship parts directly or to consolidate them through intermediate platforms.
To allow for optimized decision-making in this context, we formalize the underlying Shipper Transportation Planning Problem (STPP), tailored to the needs of a high-volume industrial shipper, and develop a scalable algorithmic framework, capable of generating high-quality plans on real data from one of Europe’s largest manufacturers.
1.1 State of the art
The design and optimization of transportation networks have been long-standing challenges in operations research, especially in the context of logistics and supply chain management. The foundational modeling framework is that of multicommodity network flow problems, extensively studied since the 1960s (Ahuja et al., 1988). Early research already recognized the critical importance of flow consolidation in freight transportation (Powell and Sheffi, 1983). More recently, Crainic et al. (2021) provided a comprehensive overview of methods for solving two closely related problems: multicommodity multifacility network design and the load plan design problem. While these models are more tractable, they exhibit many of the same structural and computational challenges as our setting.
Exact solution methods for such problems typically rely on mathematical programming techniques, including decomposition methods such as Lagrangian relaxation, Benders decomposition, and Dantzig-Wolfe reformulation (Crainic and Gendron, 2020). These methods are effective on moderate-size instances, involving a few hundred nodes and commodities, but struggle to scale. Studies by Gendron and Larose (2014) and Frangioni et al. (2017) report computation times of several hours even on such modest networks.
To tackle larger instances, researchers have focused on heuristic and matheuristic methods. Metaheuristics like large neighborhood search and population-based methods have shown promising performance when optimizing realistic-sized networks (Crainic and Gendreau, 2021; Kazemzadeh et al., 2022; Gendron et al., 2018; Paraskevopoulos et al., 2016). For example, in the context of road transportation planning, Bakir et al. (2021), Erera et al. (2013) as well as Lindsey et al. (2016) have addressed networks with tens of thousands of nodes and hundreds of thousands of arcs using sophisticated local descent-based matheuristics. Recently, Eom et al. (2025) proposed a recursive partitioning and batching framework to improve scalability and computation time.
However, these existing approaches face severe limitations when being applied to shipper-side planning at the scale encountered in our context. While some algorithms have been tested on networks of comparable size, they typically address only tens of thousands of commodities, work on time horizons ranging from a day to a week and do not consider explicit bin-packing constraints. This omission is crucial: once bin-packing is included, standard mathematical programming models become intractably large, while decomposition subproblems become too complex or numerous, and existing matheuristics fail to scale. This highlights a major gap in the literature: while LPDPs and large-scale network design problems have been studied from the perspective of carriers and 3PLs, the shipper’s perspective—particularly with explicit consolidation and massive commodity volume—remains underexplored.
1.2 Contributions
This paper studies the shipper-side design of large-scale inbound transportation networks, focusing on the strategic planning problem faced by Renault. In this context, we contribute (i) a novel, scalable metaheuristic framework that advances the methodological frontier of shipper-side network design, (ii) the first computational results on instances of genuine global industrial scale, and (iii) a set of managerial guidelines that directly inform decision-making in large-scale inbound logistics.
Methodologically, we develop a novel Iterated Local Search (ILS) metaheuristic tailored to the Shipper Transportation Design Problem. The approach integrates efficient large-neighborhood search operators with a perturbation scheme based on tractable MILP relaxations. A key innovation lies in exploiting bundle-specific decompositions and giant container relaxations to obtain scalable lower bounds, which are further used to design a rounding heuristic benchmark. To our knowledge, this is the first approach that combines these methodological ingredients into a unified framework capable of handling networks of industrial scale.
Empirically, we demonstrate that the proposed ILS achieves substantial performance improvements over both Renault’s current planning solutions and a suite of established benchmark heuristics. On the largest “World” instances, the algorithm closes the gap to the best available lower bound to within 7.9% and delivers solutions that reveals a 23.2% potential for Renault’s operational plans cost reduction. Importantly, this represents the first computational study that solves a shipper-side transportation design problem at a scale solving instances with more than 700,000 commodities and 1,200,000 arcs.
Beyond methodological advances, our analysis provides actionable managerial insights. First, we show that a faithful representation of bin-packing consolidation is indispensable for realistic network design, as simplified capacity approximations lead to substantial distortions. Second, we highlight the role of transport regularity: highly regular plans not only reduce computational complexity but also align with planners’ preference for operational stability, while more flexible plans yield only marginal additional cost savings. Third, we analyze outsourcing decisions and reveal a clear volume-dependent threshold: outsourcing may be attractive in small or low-volume networks, but for global, high-volume supply chains such as Renault’s, in-house planning and execution consistently outperform.
2 Case study & Problem Setting
This section introduces the empirical and operational context underlying the transportation planning problem addressed in this paper. We focus on the inbound supply network of Renault, a global automotive manufacturer operating a high-volume, multi-tier logistics system. The aim is to characterize the network topology, flow structure, and planning requirements that shape the problem formulation developed in the subsequent sections. We begin by describing the global supply network’s physical architecture, highlighting its hub-and-spoke structure and consolidation logic. We then analyze key flow characteristics—volume patterns, shipment granularity, and cost distributions—to identify the core logistical challenges. Finally, we present an informal problem statement that synthesizes these elements into a shipper-side transportation planning problem, capturing the scale, constraints, and cost trade-offs specific to this context.
2.1 Supply Network
Figure 1 illustrates the structure of Renault’s supply network. It includes all suppliers and industrial sites involved in the transport of car parts, as well as all logistics platforms used to consolidate flows.
(a) Suppliers
(b) Platforms
(c) Industrial Sites
Figure 1: Worldwide Renault’s sites
These maps highlight the global scale and complexity of Renault’s supply network, which comprises more than 3,000 sites, 40 industrial sites, and 100 logistics platforms. The term platforms refers to any type of logistics facility where material flows can be consolidated or unconsolidated as needed. The network is predominantly concentrated in Europe, which accounts for 70% of all sites, and consists mainly of suppliers, which account for 90% of all sites in the network. Operating at this scale introduces substantial logistical challenges. Lead times can vary from days to several weeks, depending on the route and transport mode. Intercontinental coordination, sensitivity to global disruptions (e.g., port delays, geopolitical events), and the need to synchronize flow consolidation across multiple tiers (a long-haul core network of platforms and a vast periphery of suppliers and industrial sites), all add to the complexity of transportation planning.
The network exhibits a hub-and-spoke structure: suppliers form a sparsely connected periphery, each typically linked to only one or two other nodes, while platforms act as high-degree hubs. These platforms enable flow consolidation and serve as key transshipment points that reduce the number of required direct supplier–site connections. As a result, the core network of platforms and industrial sites, although relatively small in size, is densely connected and serves as the primary backbone for flow consolidation and routing. This hub-and-spoke structure is a common design principle in large-scale industrial supply networks. By concentrating inbound flows through logistics platforms, firms can exploit economies of scale in transportation, reduce the number of direct connections between suppliers and industrial sites, and enhance operational control. Such configurations also improve scalability and resilience, as they allow new suppliers to be integrated through local connections to existing hubs and enable rerouting in response to disruptions. In the context of global automotive manufacturing, this structure supports efficient coordination across multiple tiers, regions, and transport modes. However, fully realizing these benefits requires advanced transportation planning capabilities to manage the resulting complexity in routing, consolidation, and timing decisions.
We refer to transportation between two nodes as a leg. Figure 2 visualizes all such legs within the network.
(a) Supplier $\to$ Platform
(b) Platform $\to$ Platform
(c) Platform $\to$ Industrial Sites
Figure 2: Worldwide Renault’s legs
About 80% of them connect suppliers to industrial sites and typically last less than two weeks. Another 15%, mostly intra-continental, either originate from or terminate at platforms. The remaining legs connect platforms to each other and account for the rest of the intra-continental links and nearly all inter-continental ones, with durations ranging from three to eight weeks. These platform–platform legs often rely on long-haul modes such as deep-sea shipping and require early capacity reservation and schedule coordination. The heterogeneity in leg types has important operational implications. Supplier–site legs are typically managed within short-term tactical planning horizons, whereas platform–to-platform connections involve strategic decisions regarding routing, frequency, and transport capacity. Moreover, intercontinental legs–though less frequent–represent a disproportionately large share of transportation cost and carbon emissions. Consolidating flows at platforms is therefore essential not only for cost efficiency but also for meeting sustainability targets.
In summary, the network features a distinctive topology: a compact, high-density core that enables efficient long-haul transport and consolidation, and a vast, low-connectivity periphery that generates most of the supply variability. These structural characteristics impose specific methodological requirements on transportation planning models and motivate the modeling approach developed in the following sections.
2.2 Commodities and Flows
Renault’s inbound supply chain supports a highly complex and large-scale operation. Over a six-month horizon, the network handles more than 40,000 distinct car parts, divided into 700,000 commodities to deliver, corresponding to approximately 9,000,000 packages and 20,000,000 m3 of volume. Figure 3 illustrates the distribution of this total volume over time and across shipment sizes.
(a) Volume per Week
(b) Volume per Size
Figure 3: Commodities volume distributions
The volume of inbound flows is relatively stable over time, without pronounced peaks or downswings. This temporal regularity supports the implementation of structured mid-term planning processes. For example, it enables the use of cyclic transport plans, fixed service frequencies, and pre-booked capacities for recurrent flows. Such stability reduces the need for reactive short-term adjustments and improves predictability across the network. However, the transportation plan must remain adaptable, as adjustments may occur frequently due to changes in supplier configurations or demand patterns. Figure 4 illustrates the frequency of order and the number of commodities for origin-destination pairs, differentiating between short and long distances. In this context, the sheer number of part types introduces substantial planning complexity. The high granularity of the commodity mix leads to fragmented demand patterns, with each part potentially exhibiting different frequencies, sizes, and origin–destination characteristics. This heterogeneity increases the dimensionality of planning problems and necessitates sophisticated aggregation, routing, and consolidation strategies to efficiently organize transportation.
(a) Orders
(b) Commodities
Figure 4: Distribution of Orders and Commodities inside Origin-Destinations (OD)
Most individual shipments fall within a moderate size range of 1 to 4 cubic meters, as shown in figure (3b). While these volumes are manageable, they typically do not fill a transport unit on their own. Consequently, high utilization of containers or trucks depends on consolidating flows across multiple dimensions—such as combining different part types, suppliers, or destinations within the same shipment. This introduces additional decision layers involving timing, compatibility constraints, and routing coordination, especially for multi-stop or multi-leg routes. The current organization of flows results in nearly 500,000 transport units—trucks or shipping containers—being used every six months. Their spatial distribution is shown in Figure 5.
(a) Supplier $\to$ Platform
(b) Platform $\to$ Platform
(c) Platform $\to$ industrial sites
Figure 5: Legs used in current flows
Interestingly, only a small fraction of all potential legs in the network are actively used. This is a direct consequence of the underlying consolidation strategy: flows are channeled through a limited number of high-volume corridors, where cost efficiency can be maximized. In modeling terms, this indicates that leg activation is not static but endogenous—dependent on shipment densities, consolidation opportunities, and transport cost structures. While most of the physical volume moves within Europe, the cost distribution exhibits a contrasting pattern.
Figure 6 illustrates the volume and cost repartition between flows shipped on direct legs and others. The majority of transportation costs stem from intercontinental flows, which are fewer in number but far more expensive per unit. This cost–volume asymmetry has important implications for planning. Intercontinental shipments require long lead times, early capacity reservation, and coordination with deep-sea or multimodal transport services. In contrast, intra-European flows dominate operational complexity due to their higher frequency, network density, and time sensitivity. Intercontinental flows are typically routed through designated global platforms that serve as strategic hubs for mode transitions and consolidation. These hubs enable coordination across continents and help align long-haul shipments with downstream distribution schedules in Europe. Their positioning and function are critical to synchronizing the global and regional components of the supply chain and mitigating the risk of disruptions or bottlenecks.
(a) Volume repartition
(b) Cost repartition
Figure 6: Repartition of volumes and cost between direct and non-direct shipments
In summary, the flow structure in Renault’s inbound network exhibits several defining characteristics: high commodity granularity, stable temporal distribution, uneven cost–volume structure, and selective leg utilization. These characteristics collectively impose a complex set of requirements on transportation planning models. Effective solutions must operate across multiple temporal and spatial scales, leverage consolidation potential, and dynamically determine which legs and routes to activate.
2.3 Planning problem
As stated in the introduction, Renault’s transportation planning process aims to generate a provisional six-month plan for delivering car parts from suppliers to assembly industrial sites across its global logistics network. This problem arises directly from the structural and operational characteristics discussed in the previous sections: a hub-and-spoke network with a dense platform core, high-volume and fragmented part flows, and a pronounced asymmetry between cost and volume distributions. The resulting task belongs to the class of Shipper Transportation Planning Problems, which we may describe informally as follows:
[…]
At its core, the planning problem involves routing a large number of distinct commodities through a multi-tier logistics network while satisfying operational constraints. Each commodity is defined by its origin (supplier), destination (industrial site), and required delivery window. The objective function reflects the multi-dimensional cost structure observed in Renault’s inbound system: it includes per-unit transport costs (e.g., for trucks or containers, including carbon emissions cost), volume-based handling costs at consolidation platforms, and distance-based costs associated with inventory holding or capital-in-transit. These costs capture the trade-offs between direct point-to-point shipments and the use of intermediate platforms for consolidation—an essential aspect given the network’s limited number of active legs and the prevalence of medium-sized shipments.
The constraints reflect the operational realities introduced earlier. Flows must be feasible in terms of available routes, transport modes, and delivery timing. Delivery times are modeled flexibly, allowing for bounded lead-time variation to reflect typical transport delays and scheduling options. Regularity constraints promote temporal stability in the choice of paths—motivated by the observed preference for predictable routing patterns in high-frequency intra-European flows. Explicit bin-packing constraints are necessary to model the consolidation of multiple parts into transport units. This modeling choice is critical for Renault for two reasons. First, the company has detailed knowledge of product packaging and dimensional properties, which enables exact packing representations. Second, the diversity of part volumes—ranging from small components to larger assemblies—makes simple approximations, such as integer multiples of trailers, insufficient for high-quality planning.
In summary, the planning problem reflects the need to manage a large-scale, highly fragmented flow network with variable lead times, consolidation constraints, and a mix of tactical and strategic cost drivers. The formulation integrates structural insights from Renault’s supply network with operational constraints derived from its commodity and flow characteristics. The remainder of this paper focuses on developing a scalable, high-quality solution approach for this problem.
3 Encoding structure and regularity through time expanded graphs
The transportation planning problem faced by Renault involves a highly complex decision space due to the scale of operations and the rich structure of operational constraints. Each planning instance involves millions of part-level shipments over a multi-tier network, with temporal dependencies and consolidation requirements that span across several time periods. In addition to traditional network flow constraints, the problem introduces explicit bin-packing requirements for transport units, which substantially increase the computational complexity. This packing structure not only determines the feasibility of flow assignments, but also couples them across commodities and arcs, resulting in a combinatorial explosion of the solution space. To address these challenges, we develop a graph-based modeling framework that integrates network structure, temporal dynamics, and bin-level consolidation decisions in a scalable and modular way. Beyond, providing a formal problem formulation, the structure of this modeling framework eases the development of a corresponding algorithm in the subsequent section.
Specifically, we present a formal model composed of four main components: (i) the static supply network and the temporal structure of commodities in Section 3.1, (ii) two coupled time-expanded graphs to represent delivery timing and regularity in Section 3.2, (iii) the flow, regularity, and bin-packing constraints in Section 3.3, and (iv) the associated cost structure and resulting mixed-integer programming (MIP) formulation in Section 3.4. We conclude with a discussion of model scalability and possible extensions.
3.1 Modeling the Supply Network and Commodities
We model the supply network as a directed graph $D=(V,A)$, where the node set $V=S\cup P\cup U$ includes all physical locations in the network. The set $S$ contains suppliers, $P$ represents logistics platforms, and $U$ denotes production units. The arc set $A\subseteq V^{2}$ captures all possible transportation links between locations and is partitioned into four categories. Collection arcs $A^{\mathrm{col}}$ connect suppliers to platforms and represent the first leg of multi-leg paths. Inter-platform arcs $A^{\mathrm{pla}}$ connect different platforms and enable the transfer of consolidated goods within the intermediate logistics network. Delivery arcs $A^{\mathrm{del}}$ link platforms to production units and represent the final leg of consolidated flows. Finally, direct arcs $A^{\mathrm{dir}}$ connect suppliers directly to production units, bypassing intermediate consolidation. Figure 7 illustrates a representative example of such a network, comprising two suppliers, three platforms, and one production unit.
Figure 7: Example of network graph $D$
The inbound supply chain operates on a rolling time horizon $T]={1,\ldots,T}$, discretized into weekly steps. A commodity $m\in M$ represents a non-splittable part sent from supplier $s_{m}\in S$ to unit $u_{m}\in U$ for delivery in week $d_{m}\in[T]$. Each commodity has a volume $\ell_{m}\in\mathds{R}$, a maximum delivery time $\tau_{m}\in\mathds{N}$ (in weeks), and a multiplicity $q_{m}$ representing the quantity to be delivered in week $d_{m}$. The specific delivery path used by each commodity is a decision variable. In this context, we note that we treat each supplier–unit–delivery tuple as a distinct commodity, although the same part may be sourced from multiple suppliers or sent to multiple industrial sites. This simplification aligns with the scope of this work, which focuses on transportation planning rather than sourcing decisions.
To structure delivery decisions and ensure operational regularity, we introduce two key constraints: grouping and time regularity. Grouping enforces that all commodities sharing the same origin supplier, destination unit, and delivery date must follow an identical path through the network. To formalize this, we aggregate such commodities into orders $o\subseteq M$. Each order $o$ is characterized by a common delivery date $d_{o}$, a shared maximum delivery time $\tau_{o}$, and a total volume $\ell_{o}=\sum_{m\in o}q_{m}\ell_{m}$. Let $O$ denote the set of all orders. Time regularity imposes consistency in routing over time: commodities with identical supplier–unit pairs must follow the same sequence of nodes in the supply network, regardless of their delivery date. To capture this, we group all orders with the same supplier and unit but differing delivery dates into bundles $b\subseteq O$, which represent recurring flows along regular paths. Let $B$ denote the set of all bundles.
Together, these two layers of aggregation form a hierarchical structure: the set of orders $O$ partitions the set of commodities $M$, and the set of bundles $B$ partitions the set of orders $O$. This structure ensures that all commodities within a bundle follow the same node sequence, thereby enforcing both intra-week consistency (via grouping) and inter-week regularity (via bundling) in delivery planning.
3.2 Time-Expanded Graph Structures
To model both the temporal flow of commodities and the regularity of delivery paths, we define two complementary time-expanded graphs. Figure 8 illustrates these two time expansions and their relations, while purposely omitting rolling-horizon forming arcs for clarity.
Figure 8: Example of time-space digraph $\mathcal{D}$ (top) and travel-time digraph $\mathscr{D}$ (bottom)
Note : Given a bundle $b\in B$, we compute the common path $\phi_{b}$ used by all orders $o\in O^{b}$ on the travel-time graph $\mathscr{D}$ (in red on the bottom graph). Given an order $o\in b$ we translates this common path with $\pi_{\mathcal{D}}$ into a path $\varphi_{o}=\pi_{\mathcal{D}}(\phi_{b},o)$ on $\mathcal{D}$ (in red on the top graph), removing shortcut arcs used and recovering the actual timed path used by the commodities $m\in o$.
Time-Space Graph
To capture the temporal dynamics of the supply network, we expand the static graph $D$ into a time-indexed structure that reflects both fluctuating demand and transportation durations over the planning horizon. This results in the time-space graph, denoted by $\mathcal{D}=(\mathcal{V},\mathcal{A})$. We define the set of timed nodes as $\mathcal{V}=V\times[T]$, where each node $\nu=(v,t)$ represents location $v\in V$ at time step $t\in[T]$. For each arc $a=(u,v)\in A$, we let $\tau_{a}\in\mathds{N}$ denote the number of time steps required to traverse the arc. Using this, we define the set of timed arcs as
$\mathcal{A}=\left\{\big{(}(u,t),(v,t^{\prime})\big{)}\;\middle|\;(u,v)\in A\;\text{and}\;t^{\prime}\equiv t+\tau_{(u,v)}\pmod{T}\right\}\subseteq\mathcal{V}^{2}$
The modular operator rolls the time horizon, allowing arcs that start near the end of the horizon to wrap around to the beginning of the next cycle. We use the symbol $a$ for arcs in either $A$ or $\mathcal{A}$, relying on context to distinguish between the two.
Travel-Time Graph
To model regularity constraints and account for flexible delivery times, we define the travel-time graph $\mathscr{D}=(\mathscr{V},\mathscr{A})$. This graph is a partial time-expansion of the network graph $D$, where time steps represent the remaining time until delivery rather than absolute positions on the planning horizon. Unlike the time-space graph, which expands over the full horizon $[T]$, the travel-time graph expands over the partial horizon $\mathcal{T}]$, where $\mathcal{T}=\max_{o\in O}\tau_{o}$ denotes the maximum delivery time allowed for any order. We expand suppliers and platforms across all time steps in $\mathcal{T}]$, while production units only appear at the final time step $\mathcal{T}$. Accordingly, we define the node set as $\mathscr{V}=\mathscr{S}\cup\mathscr{P}\cup\mathscr{U}$, where $\mathscr{S}=S\times[\mathcal{T}]$, $\mathscr{P}=P\times[\mathcal{T}]$, and $\mathscr{U}=U\times\{\mathcal{T}\}$. We construct the arc set $\mathscr{A}$ analogously to that of the time-space graph, with one key extension: we introduce shortcut arcs to allow delivery time flexibility. These arcs are defined as $\alpha\in\mathscr{A}^{\mathrm{short}}=\{((s,t),(s,t-1))\mid(s,t)\in\mathscr{S}\}$ and represent the option to delay the dispatch of a commodity by one time unit at its origin.
Interactions
The travel-time graph $\mathscr{D}$ interacts with the time-space graph $\mathcal{D}$ through a projection operator $\pi_{\mathcal{D}}$, which maps nodes in $\mathscr{D}$ to corresponding nodes in $\mathcal{D}$. For any node $\upsilon=(v,t)\in\mathscr{V}$ in the travel-time graph, the index $t$ represents the number of time steps remaining until delivery. Given an order $o\in O$ with delivery date $d_{o}$, the corresponding node in the time-space graph is $\nu=\pi_{\mathcal{D}}(\upsilon,o)=(v,d_{o}-t)\in\mathcal{V}$. This projection mechanism enables a consistent translation between the two graph representations. While the travel-time graph serves to compute bundle paths that satisfy regularity constraints and flexible delivery times, the time-space graph tracks the resulting flows over the rolling planning horizon, incorporating time-dependent demand and transport durations. Although maintaining and synchronizing both graphs introduces additional modeling complexity, it allows the majority of path computations to occur on $\mathscr{D}$, which is significantly smaller than $\mathcal{D}$, thus improving computational efficiency.
3.3 Flow, Regularity, and Bin-packing constraints
Flows and Regularity Constraints
We denote by $f_{a}^{m}\in\mathds{N}$ the quantity of commodity $m\in M$ flowing on arc $a\in\mathcal{A}$. We denote by $x_{\alpha}^{b}\in\{0,1\}$ the binary variable indicating whether bundle $b\in B$ uses arc $\alpha\in\mathscr{A}$. The flow and regularity constraints define an elementary path on the travel-time graph, as captured by equations (2) and (3), and project the corresponding quantities onto the time-space graph, as shown in equation (4). For $b\in B$ and $\nu\in\mathscr{V}$, $e_{\nu}^{b}=1$ for $\nu=(s_{b},\mathcal{T}-\tau_{b})$, 1-1 for $\nu=(u_{b},\mathcal{T})$ and 0 otherwise.
$\sum_{\alpha\in\delta^{+}(\nu)}x_{\alpha}^{b}-\sum_{\alpha\in\delta^{-}(\nu)}x_{\alpha}^{b}=e_{\nu}^{b}\quad\forall b\in B\text{ , }\nu\in\mathscr{V}$ (2)
$\sum_{t\in[\mathcal{T}]}\sum_{\alpha\in\delta^{-}((p,t))}x_{\alpha}^{b}\leq 1\quad\forall b\in B\text{ , }p\in P$ (3)
$f_{a=\pi_{\mathcal{D}}(\alpha,o)}^{m}=q_{m}x_{\alpha}^{b}\quad\forall b\in B\text{ , }o\in b\text{ , }m\in o\text{ , }\alpha\in\mathscr{A}$ (4)
The elementarity constraint is needed because relaxing it leads to paths that go through the same node at different time steps as such paths enable to improve consolidation.
Transportation and Bin Packing
Transport on arcs $a\in\mathcal{A}$ involves consolidation. The shipper procures transport units, referred to generically as bins, which typically correspond to trucks or ship containers. All commodities assigned to a given arc must be packed into these bins, requiring the solution of a bin-packing problem. In practice, this problem is multi-dimensional and subject to additional constraints. However, for tractability, we approximate it using a classical single-dimensional bin-packing formulation. We assume that each bin is fully loaded at the origin and completely unloaded at the destination of its respective arc.
For each arc $a\in\mathcal{A}$, let $L_{a}\in\mathds{R}$ denote the bin capacity and $K_{a}\in\mathds{N}$ the maximum number of bins available. We introduce decision variables $\tau_{a}^{k}\in\{0,1\}$ to indicate whether bin $k\in[K_{a}]$ is used, and $y_{ak}^{m}\in\mathds{N}$ to represent the number of units of commodity $m$ assigned to bin $k$. This bin-packing model yields the constraints (5) and (6).
$f_{a}^{m}=\sum_{k\in K}y_{ak}^{m}\quad\forall m\in M\text{ , }a\in\mathcal{A}^{\mathrm{con}}$ (5)
$\sum_{m\in M}y_{ak}^{m}\ell_{m}\leq L_{a}\tau_{a}^{k}\quad\forall a\in\mathcal{A}^{\mathrm{con}}\text{ , }k\in K_{a}$ (6)
Outsourcing Exception
Transportation on a subset of arcs, specifically $a\in\mathcal{A}^{\mathrm{out}}\subset\mathcal{A}$, is outsourced. In this case, the shipper procures transport services from a third-party logistics provider (3PL) on a per-commodity basis, rather than contracting directly with a carrier. As a result, there is no need to explicitly model bin-level consolidation on these arcs. To distinguish between outsourced and consolidated flows, we define the set of consolidated arcs $\mathcal{A}^{\mathrm{con}}\subseteq\mathcal{A}$, which comprises all arcs requiring explicit bin-packing and consolidation. We therefore have $\mathcal{A}^{\mathrm{out}}\;\uplus\;\mathcal{A}^{\mathrm{con}}=\mathcal{A}$. In Renault’s case, we even have $\mathcal{A}^{\mathrm{out}}\subseteq\mathcal{A}^{\mathrm{col}}$.
3.4 Cost Structure and Full Problem Formulation
This subsection presents the complete cost structure and optimization formulation for the shipper transportation planning problem.
Transportation
The model assigns a transport cost for each bin used on consolidated arcs $a\in A^{\mathrm{con}}$, denoted by $c_{a}^{\mathrm{con}}$. On outsourced arcs $a\notin A^{\mathrm{con}}$, the shipper incurs a volume-based cost $c_{a}^{\mathrm{out}}$. Each arc also incurs a volume-dependent cost $c_{a}^{\mathrm{CO2}}$ based on its carbon emission factor.
Platforms
Processing and handling goods at logistics platforms generates costs. The platform cost $c_{p}^{\mathrm{plat}}$ increases proportionally with the volume processed. Platform contracts define a capacity limit $u_{p}^{\mathrm{plat}}$, specifying the maximum volume allowed at the base rate. Exceeding this capacity leads to an overload cost $c_{p}^{\mathrm{over}}$. Let $z_{p}\in\mathds{R}$ represent the excess volume at platform $p$. The model enforces the following overload constraint:
$\sum_{a\in\delta^{-}(p)}\sum_{m\in M}f_{a}^{m}\ell_{m}\leq u_{p}^{\mathrm{pla}}+z_{p}\quad\forall p\in\mathcal{P}$ (7)
Commodities
Commodities in transit tie up capital, which we model using a capital cost $c_{m}^{\mathrm{cap}}$, proportional to the distance traveled. The cost incurred by commodity $m$ on arc $a$ is given by $c_{am}^{\mathrm{cap}}$.
$c_{am}^{\mathrm{cap}}=d_{a}\times c_{m}^{\mathrm{cap}}$
Network Cost Function
To simplify notation, we aggregate all commodity-specific costs for arc $a=(u,v)$ into a composite term $c_{am}^{\mathrm{com}}$.
$c_{am}^{\mathrm{com}}=\frac{\ell_{m}}{L_{a}}c_{a}^{\mathrm{CO2}}+c_{am}^{\mathrm{cap}}+\mathds{I}(v\in\mathcal{P})\frac{\ell_{m}}{L_{a}}c_{v}^{\mathrm{plat}}+\mathds{I}(a\notin\mathcal{A}^{\mathrm{con}})\frac{\ell_{m}}{L_{a}}c_{a}^{\mathrm{out}}$
Here, $\mathds{I}(v\in\mathcal{P})$ equals 1 if node $v$ is a platform, and 0 otherwise. Likewise, $\mathds{I}(a\notin\mathcal{A}^{\mathrm{con}})$ equals 1 if arc $a$ is outsourced. Then, the total network cost is:
$\sum_{a\in\mathcal{A}}\left(\sum_{k\in K}c_{a}^{\mathrm{con}}\tau_{a}^{k}+\sum_{m\in M}c_{am}^{\mathrm{com}}f_{a}^{m}\right)+\sum_{p\in\mathcal{P}}c_{p}^{\mathrm{over}}z_{p}$ (8)
This objective captures the cost of using bins on consolidated arcs, volume-dependent costs on all arcs, and overload penalties at platforms.
Shipper Transportation Planning Problem
The complete Shipper Transportation Planning Problem can now be formulated as the following mixed-integer program:
[…]
$\text{Flow variables }x\in\{0,1\},\ f\in\mathds{N}$
$\text{Packing variables }\tau\in\{0,1\},\ y\in\mathds{N}$
$\text{Platform variables }z\in\mathds{R}_{+}$
Even with size-reduction techniques, solving industrial-scale instances remains intractable. Typical instances feature approximately $|\mathcal{A}^{\mathrm{con}}||K||M|\approx 10^{12}$ variables and $|B||\mathscr{V}|\approx 10^{10}$ constraints, far exceeding the capability of current (commercial) solvers.
Discussion and Extensions
Although tailored to Renault’s operations, the model accommodates a wide range of extensions, including multi-modal transport, vehicle tours, inventory, part sourcing and return logistics. Appendix A provides details for implementing these extensions. While the remainder of this paper focuses on Renault’s specific operational setting, the proposed model and the solution approach developed accommodates with ease the additional features, enabling their usage to a broad range of planning environments.
4 Decomposition-guided ILS
The problem instances under consideration involve far too many variables and constraints, rendering exact methods such as mixed-integer programming or classical decomposition techniques computationally infeasible. To address this challenge, we develop a tailored Iterated Local Search (ILS) algorithm designed for large-scale combinatorial optimization under structural constraints. Described in Figure 9, it works as follows.
Figure 9: Description of the Iterated Local Search
Note : A constructive heuristic followed by a local search finds a promising candidate solution. The search space is then explored and the solution refined by alternating between a local search and a perturbation phase.
Given an instance $I$, a constructive heuristic first builds an initial solution $S$, which is then refined by alternating between a local search that improves the current solution locally and a perturbation phase that gets out of local minima.
The design of the ILS algorithm is guided by four key observations: i) the problem becomes tractable when restricted to a single commodity bundle; ii) Due to consolidation effects, bundles rarely follow their individual shortest paths in the network; iii) A significant portion of transportation costs is concentrated on shared network segments; iv) Bundles relying on these segments are typically intercontinental or low-volume, making them especially dependent on consolidation.
Based on these insights, we decompose the overall algorithm and its components into three main types of computational routines: packing, single-bundle, and multi-bundle operations. Each component, defined as a variant of a routine, plays a specific role within the ILS framework—constructing initial solutions, refining existing ones, or exploring larger neighborhoods via controlled perturbation. Table 1 summarizes the computational building blocks and their respective variants, including the geometric principles they exploit and their functional roles within the algorithm.
Table 1: Computation types and their variants
Type Variants Geometric Insight Usage Description
Packing Pack Batched Bin Packing All Inserts commodities on arcs; computes associated cost
Re-Pack Classical Bin Packing Local Search Removes and re-packs all commodities to allow cost re-optimization
Single Bundle Insertion Shortest Path on Sparse Graph (weighted with Pack) Constructive Heuristic Builds paths one commodity at a time using sparse cost structure
Re-Insertion Shortest Path on Sparse Graph (weighted with Pack) Local Search Removes and reinserts bundles to escape local minima
Multi-Bundle Consolidate Shortest Path on Sparse Graph (weighted with Pack) Local Search Merges bundles on shared paths before re-insertion; partial path recomputation
Flow-based Partial MILP Relaxation Perturbation Solves aggregated routing via relaxed MILP to explore new configurations
Path-based Partial MILP Relaxation Pertubation Similar to flow-based, but works on enumerated route alternatives
4.1 Constructive algorithm
The constructive heuristic incrementally builds a feasible solution $S$ for a given instance $I$ by inserting commodities bundle by bundle. At each step, it chooses configurations that minimize arc activation and platform usage costs by solving a bundle insertion sub-problem. Figure 10 illustrates the heuristic’s logic. The heuristic follows the classical First-Fit Decreasing (FFD) rule from bin packing (Korf, 2002). FFD processes items in decreasing order of size and inserts each into the first bin with sufficient space. In our setting, the heuristic inserts bundles $b\in B$ into the partial solution $S$ in decreasing order of their maximum packaging size $\ell_{B}=\max_{m\in B}\ell_{m}$, which serves as a proxy for packing complexity. At each iteration, the heuristic inserts a bundle in the current solution using a minimum cost insertion subproblem and maintains a partial solution that satisfies constraints (2)–(6) for the current set of bundles $B^{S}\subset B$. Each bundle $b\in B^{S}$ is associated with a path $\phi_{b}$, and the set of inserted commodities grows as the algorithm progresses.
Figure 10: Description of the Constructive Heuristic
Note : Bundles are sorted by decreasing size and then inserted one by one in the current solution.
Minimum Cost Insertion
To insert a new bundle $b$ into the partial solution $S$ at minimum cost, the algorithm solves a restricted version of problem (9). This subproblem identifies a feasible delivery path for $b$ and evaluates the resulting consolidated arc costs as follows: define $M^{S}=\bigcup_{b\in B^{S}}\{m\in b\}$ as the set of commodities already in $S$, and let $(\bar{y},\bar{f})$ denote their fixed packing decisions. Using those, it enforces solution-aware packing and platform constraints when inserting $b$ : bin capacity $L_{a}$ becomes $\bar{L}_{a}^{k}$ and platform capacity $u_{p}^{\mathrm{pla}}$ becomes $\bar{u_{p}}^{\mathrm{pla}}$.
$\bar{L}_{a}^{k}=L_{a}-\sum_{m\in M^{S}}\bar{y}_{ak}^{m}\ell_{m}\quad\text{ and }\quad\bar{u_{p}}^{\mathrm{pla}}=u_{p}^{\mathrm{pla}}-\sum_{a\in\delta^{-}(p)}\sum_{m\in M^{S}}f_{a}^{m}\ell_{m}$
The insertion subproblem $\mathcal{P}_{\mathrm{Insertion}}(S,b)$ is defined as follows:
[…]
$x,\tau\in\{0,1\},\quad y,f\in\mathds{N},\quad z\in\mathds{R}_{+}$
To simplify notation, the algorithm assumes all decision variables for bundles not yet inserted are set to zero.
Solving the Minimum Cost Insertion
The algorithm solves problem (10) efficiently by reducing it to a shortest path problem on a specialized subgraph of $\mathscr{G}$. Since the subproblem considers a single bundle, the algorithm can precompute all costs and project them from $\mathcal{G}$ onto $\mathscr{G}$. The bundle-specific digraph $\mathscr{G}^{b}=(\mathscr{V}^{b},\mathscr{A}^{b})$ contains only nodes and arcs relevant to bundle $b$, defined as $\mathscr{V}^{b}=\{\nu=(v,t)\in\mathscr{V}\mid v\in\{s_{b},u_{b}\}\cup P\},\;\mathscr{A}^{b}=\{\alpha=(u,v)\in\mathscr{A}\mid u,v\in\mathscr{V}^{b}\}.$ In practice, the set $\mathscr{V}^{b}$ is further pruned by removing platforms that cannot be reached from $s_{b}$ or cannot reach $u_{b}$. For each arc $\alpha\in\mathscr{A}^{b}$, the algorithm computes three types of costs: the commodity cost $c_{\alpha}^{\mathrm{com}}$, the platform overloading cost $c_{\alpha}^{\mathrm{over}}$ if arc $\alpha$ ends at a platform $p\in P$ and finally the consolidated arc cost $c_{\alpha}^{\mathrm{con}}$. The commodity and the platform overloading costs are the sum of corresponding costs for each projected arc.
$c_{\alpha}^{\mathrm{com}}=\sum_{\begin{subarray}{c}o\in b\\ a=\pi_{\mathcal{D}}(\alpha,o)\end{subarray}}\ell_{o}\,c_{am}^{\mathrm{com}}\quad\text{ and }\quad c_{\alpha}^{\mathrm{over}}=\sum_{\begin{subarray}{c}o\in b\\ a=\pi_{\mathcal{D}}(\alpha,o)\end{subarray}}c_{p}^{\mathrm{over}}\cdot\max\left(0,\ell_{o}+\bar{L}_{a}-u_{p}^{\mathrm{pla}}\right)$
The consolidated arc cost is computed with a bin-packing for each projected arc.
$c_{\alpha}^{\mathrm{con}}=\sum_{\begin{subarray}{c}o\in b\\ a=\pi_{\mathcal{D}}(\alpha,o)\end{subarray}}c_{a,o}^{\mathrm{con}}\quad\text{ and }\quad c_{a,o}^{\mathrm{con}}=\left\{\begin{aligned} \min\quad&\sum_{k\in K}c_{a}\tau_{a}^{k}\\ \text{s.t.}\quad&q_{m}=\sum_{k\in K}y_{ak}^{m}&&\forall m\in o\\ &\bar{L}_{a}^{k}+\sum_{m\in o}y_{ak}^{m}\ell_{m}\leq L_{a}\tau_{a}^{k}&&\forall k\in K_{a}\\ &\tau\in\{0,1\},\quad y\in\mathds{N}\end{aligned}\right$.
Each arc $\alpha\in\mathscr{A}^{b}$ is then weighted by the sum if these three components $c_{\alpha}=c_{\alpha}^{\mathrm{com}}+c_{\alpha}^{\mathrm{over}}+c_{\alpha}^{con}$
Proposition 1.
Solving the minimum cost insertion problem $\mathcal{P}_{\mathrm{Insertion}}(b)$ (10) amounts to computing an elementary shortest path from $(s_{b},\mathcal{T}-\tau_{b})$ to $(u_{b},\mathcal{T})$ in $\mathscr{G}^{b}$.
In practice, this result leads to two important implementation choices. First, the elementarity constraint (3) is nearly always inactive. In over 99% of cases, the bundle path is already elementary, allowing the use of Dijkstra’s algorithm with non-negative arc costs to efficiently solve the problem. Second, in the rare cases where elementarity must be enforced, the algorithm insert the bundle in an empty solution instead of the current solution, which guarantees that the path is going to be elementary.
To maintain tractability across the large number of arc-level bin-packing problems—up to $O(10^{5})$ instances—the algorithm applies the FirstFitDecreasing heuristic. This approach ensures near-optimal performance in practice while keeping computational effort manageable. It also satisfies a known approximation guarantee (Dósa, 2007): $\forall I,\ \text{FFD}(I)\leq\frac{11}{9}\cdot\text{OPT}(I)+\frac{6}{9}.$
4.2 Local Search
The local search algorithm operates through three custom-designed and increasingly large neighborhoods: Re-Pack, Re-Insert, and Consolidate-and-Refine. At each iteration, it randomly selects one of these neighborhoods and generates a neighboring solution. If this neighbor improves upon the current solution, the algorithm accepts it as the new incumbent; otherwise, it discards the candidate and continues the search. Figure 11 illustrates this procedure.
The neighborhood designs draw on empirical observations from our numerical experiments. While local search methods often benefit from well-crafted neighborhoods, Turkeš et al. (2021) show that exhaustive tuning of neighborhood mechanics rarely yields significant gains. We therefore prioritized simplicity and computational efficiency. Instead of exhaustively evaluating all neighbors, the algorithm samples a single random neighbor in each iteration.
Figure 11: Description of the Local Search procedure
Note : Neighborhoods are uniformly sampled to produce candidate solutions, which becomes the current solution in case of improvement.
Re-Pack Neighborhood
In this neighborhood, the algorithm recomputes bin-packings across a set of arcs $\mathcal{A}$ given using a portfolio of heuristics $(BP_{i})_{i\in[n]}$. In practice, $\mathcal{A}$ is usually either all consolidated arcs or all projections of specific bundle paths and two heuristics are used : First-Fit Decreasing and Best-Fit Decreasing. The set of arcs is also pruned using bin-packing lower bounds. By globally recomputing packings, the algorithm avoids suboptimal local packing sequences that arise from purely sequential insertions. Figure 12 visualizes this mechanism. Since this operation cannot increase the cost of the current solution, it is always accepted when feasible.
Figure 12: Description of the Re-Pack Neighborhood
Note : Given a set of arcs $\mathcal{A}$, new bins are computed based on provided bin-packing heuristics.
Re-Insert Neighborhood
This neighborhood removes a randomly selected bundle from the current solution, adapts the affected bins and reinserts it using the insertion procedure $\mathcal{P}_{Insertion}(S,b)$ defined in (10). Figure 13 shows the process.
Figure 13: Description of the Re-Insert Neighborhood
Note : Building on the constructive heuristic, a bundle is removed and re-inserted in the adapted solution.
Consolidate-and-Refine
The third neighborhood, illustrated in Figure 14, targets consolidation commodities along the common network connecting platforms and industrial sites. It selects two nodes $u$ and $v$ and collects the set of bundles $B^{uv}$ that travel between them. It then aggregates these bundles into a temporary super-bundle $\beta=\bigcup_{b\in B^{uv}}b$ and inserts it as a single entity using the ReInsert operator. After this coarse consolidation, the algorithm re-evaluates each bundle in $B^{uv}$ individually via further ReInsert operations to refine the solution. This ensures that only bundles benefiting from consolidation remain on the new path.
Figure 14: Description of the Consolidate Neighborhood
Note : This neighborhood seeks consolidation for bundles flowing from $u$ to $v$ on this part of their paths.
4.3 Perturbation Scheme
The final component of our Iterated Local Search metaheuristic is a perturbation scheme. This module acts as a large-neighborhood search mechanism based on integer programming. It enables the algorithm to escape local minima by re-optimizing parts of the solution and exploring promising regions of the search space. We refer to these large-scale modifications as perturbations because they rely on approximate subproblems. Although this approximation increases tractability, it also means that solutions generated through perturbation may not always improve upon the incumbent. A cost increase tolerance is therefore set 1.5% to discard perturbations that would degrade the solution quality too much. Figure (15) illustrates the perturbation process.
Figure 15: Description of the Perturbation Scheme
Note : Flow-based and Path-based perturbations are alternated until the perturbation threshold is reached.
Perturbation Design
Each perturbation re-optimizes a subset of bundles $\bar{B}\subseteq B$, following a principle similar to the ReInsert operator, but on a larger scale. The algorithm removes all bundles in $\bar{B}$ from the current solution and fixes the configuration of all other bundles in $B\setminus\bar{B}$. It then solves a simplified MILP to reinsert $\bar{B}$.
This simplification replaces the detailed packing constraints (5) and (6) with a giant container approximation:
$\bar{L}_{a}+\sum_{m\in\bar{B}}f_{a}^{m}\ell_{m}\leq L_{a}\tau_{a},\quad\tau_{a}\in\mathds{N},\quad\forall a\in\mathcal{A}^{\mathrm{con}}$ (11)
To mitigate the effect of this relaxation, we scale the transport cost on arcs $a\in\mathcal{A}^{\mathrm{con}}$ using a slope-scaling heuristic inspired by Jarrah et al. (2009). The adjusted arc cost becomes $\bar{c}_{a}$.
$\bar{c}_{a}=c_{a}\cdot\frac{BP(a)}{\lceil\sum_{m\in M}f_{a}^{m}\ell_{m}\rceil}$
Here, $BP(a)$ equals the number of bins on arc $a$ in the current solution (before reoptimization).
The resulting objective function is given by:
$\sum_{a\in\mathcal{A}}\left(\bar{c}_{a}\tau_{a}+\sum_{m\in M}c_{am}^{\mathrm{com}}f_{a}^{m}\right)+\sum_{p\in\mathcal{P}}c_{p}^{\mathrm{over}}z_{p}$ (12)
We consider two formulations for implementing perturbations: flow-based reoptimization and path-based reoptimization. The former allows complete routing flexibility but is tractable only for small $\bar{B}$; the latter restricts routing to a small number of preselected paths per bundle, allowing it to scale to larger $\bar{B}$.
Flow-Based Reoptimization
This approach reinserts a subset of bundles $\bar{B}$ by solving the following MILP:
[…]
$x\in\{0,1\},\quad f,\tau\in\mathds{N},\quad z\in\mathds{R}_{+}$
We define three families of perturbations within this framework, each based on a shared bundle characteristic (Table 2). To maximize the effectiveness of this scheme, the algorithm stacks multiple subproblems until the aggregated instance reaches a variable budget that can be solved within a few minutes (on the order of several million variables). Perturbations continue iteratively until the number of modified paths exceeds a specified threshold.
Table 2: Three families of flow-based reoptimization
Name Family of Subproblem Subproblem $\bar{B}$
Single Plant industrial site $u\in U$ $\{b\in B\mid u_{b}=u\}$
Single Supplier Supplier $s\in S$ $\{b\in B\mid s_{b}=s\}$
Random Number $n\geq 1$ Uniform sample of size $n$ from $B$
Path-Based Reoptimization
This alternative formulation improves scalability for larger $\bar{B}$. Instead of computing paths during optimization, it assumes a small set of feasible paths $\varphi\in\Phi^{b}$ is given for each bundle $b\in\bar{B}$. The flow constraints are replaced by path-based formulations:
$\sum_{\varphi\in\Phi^{b}}x_{\varphi}^{b}=1,\quad x_{\varphi}^{b}\in\{0,1\},\quad\forall b\in\bar{B}$ (14)
$f_{\pi_{\mathcal{D}}(\alpha,o)}^{m}=q_{m}\sum_{\varphi\ni\alpha}x_{\varphi}^{b},\quad\forall(b,o,m)\in\bar{B},\;\alpha\in\mathscr{A}$ (15)
We define three families of path-based perturbations (Table 3), two of them adapted from strategies proposed by Lindsey et al. (2016). The Attract perturbation introduces bundles to arcs they currently avoid, while Reduce attempts to reroute bundles away from congested arcs. Finally, the Directs perturbation is used as the last perturbation before activating the local search as it tends to switch bundles on directs paths to the shared network.
Table 3: Three families of path-based reoptimization
Name Family of Subproblem Subproblem $\bar{B}$
Attract Arc $a\in\mathscr{A}^{\mathrm{pla}}\cup\mathscr{A}^{\mathrm{del}}$ $\{b\in B\mid a\notin p_{b}\}$
Reduce Arc $a\in\mathscr{A}^{\mathrm{pla}}\cup\mathscr{A}^{\mathrm{del}}$ $\{b\in B\mid a\in p_{b}\}$
Directs $\emptyset$ $\{b\in B\mid b\text{ on direct path}\}$
As in the flow-based case, we stack multiple subproblems to scale the perturbation. Since not all bundles are eligible for rerouting through a given arc, we prioritize arcs where a large share of bundles are candidates—i.e., bundles for which the arc is either currently unused (Attract) or already used (Reduce). While these MILPs are typically faster to solve, they require extensive path enumeration, which can offset their computational advantages at the scale of a full perturbation.
5 Tractable Lower Bounds
To evaluate the quality of the proposed heuristic, we consider three lower bounding procedures that offer different trade-offs between tightness and computational complexity. These procedures, summarized in Figure 16, correspond to successive relaxations of the bin-packing structure: a linear relaxation, a mixed giant container relaxation, and a full giant container relaxation. The linear relaxation can be decomposed into independent subproblems per bundle, making it computationally tractable but relatively weak. The mixed giant container relaxation strengthens this bound on direct arcs, offering a more favorable balance between quality and efficiency. Extending the giant container relaxation to all arcs yields the full giant container bound, which yields a stronger bound at the expense of increased computational costs. Finally, from the mixed giant container relaxation we derive a rounding heuristic that produces feasible solutions and serves as a benchmark for assessing the performance of our ILS algorithm.
Figure 16: Example of linear relaxation (top) and giant container relaxation (middle) for bin packing (bottom)
Linear Relaxation Bound.
We first consider the linear relaxation of problem (9). In this relaxation, the packing variables $\tau$ and $y$ as well as the packing constraints (6) and (5) can be removed. These elements serve only to determine the number of transport units used by commodities on each arc. In the relaxed setting, this number equals the total commodity volume on the arc divided by the unit capacity. We therefore obtain the following linear relaxation :
$\min$ $\sum_{a\in\mathcal{A}}\sum_{m\in M}(c_{am}^{\mathrm{com}}+\frac{\ell_{m}}{L_{a}}c_{a}^{\mathrm{con}})f_{a}^{m}+\sum_{p\in\mathcal{P}}c_{p}^{\mathrm{over}}z_{p}$ (16)
[…]
Although this formulation still involves $O(10^{10})$ variables and cannot be solved directly with current linear programming technology, it can be decomposed by bundle if the platform overloading variables $z$ and constraints (7) are omitted. This leads to the following formulation:
[…]
Proposition 2.
The linear relaxation of  (9) has the same value as problem (16). Without platform overloading constraints (7), it is equal to the sum of bundle specific subproblems (17).
Each of these bundle-specific subproblems corresponds to the linear relaxation of the minimum-cost insertion problem (10) without platform overloading constraints. The same solution approach therefore applies: the problem reduces to a shortest path computation on the bundle-specific digraph $\mathscr{G}^{b}$. For the proof of Proposition 2 we refer to Appendix B.
Mixed-Giant Container Bound.
The quality of the previous bound can be improved by exploiting a structural property of the problem. On direct arcs, only commodities from a single bundle are present, meaning that there is no coupling across bundles. For these arcs, we can therefore strengthen the relaxation by applying a giant container approximation. The arc cost then becomes
$\forall a\in\mathcal{A}^{dir}\ :\ \left\lceil\sum_{m\in b}\frac{f_{a}^{m}\ell_{m}}{L_{a}}\right\rceil c_{a}^{\mathrm{con}}+\sum_{m\in b}c_{am}^{\mathrm{com}}f_{a}^{m}$
The non-linearity introduced in the objective function can be handled in practice by precomputing the integer number of transport units required by each order and using this value to evaluate the direct arc costs in the travel-time graph.
Rounding Heuristic.
Both the linear relaxation and the mixed giant container bound produce shortest paths for each bundle. By fixing these paths and subsequently solving a bin-packing problem for every arc of the time–space graph, we obtain feasible packings and thus a valid solution to problem (9). We apply this approach using the mixed giant container bound, as it is stronger than the linear relaxation. The resulting feasible solution serves as a benchmark for our ILS algorithm.
Full Giant Container Bound.
Applying the giant container relaxation to all arcs of the time–space graph yields problem (13) without cost scaling, where the subset of bundles is $\bar{B}=B$. This formulation inspired the perturbation MILP introduced earlier. Unlike the previous bounds, the full giant container relaxation is no longer decomposable by bundle, making it tractable only for instances of moderate size. It is closely related to the load plan design problem studied in the literature (Erera et al., 2013; Lindsey et al., 2016).
6 Computational Study
To assess the practical relevance of our approach, we conduct a comprehensive computational study on real-world instances from Renault’s inbound logistics network. The analysis proceeds in four steps. We first describe the experimental design, including the instances, benchmarks, and implementation details (Section 6.1). We then examine the role of key hyperparameters and time limits in shaping the behavior of our ILS algorithm (Section 6.2). Next, we compare solution quality and runtime against alternative strategies to evaluate relative performance (Section 6.3). Finally, we extend the analysis to highlight managerial insights on consolidation, regularity, and outsourcing, providing actionable guidance for industrial practice (Section 6.4).
6.1 Experimental design
To evaluate the performance of our ILS metaheuristic, we address two main questions: (1) How do the key difficulties identified in the case study translate into the numerical resolution of real industrial instances? (2) How does our algorithm perform compared to established resolution strategies? Before answering these questions, we introduce the test instances, describe the implementation and hyperparameter choices, and present the benchmark heuristics used for comparison.
Instances
We rely on two sets of instances derived from real industrial data from Renault’s inbound logistics operations. The first set of instances, illustrated in Figure 17, is designed to address the scalability question. It consists of five instances of increasing size: Small (S), Medium (M), Large (L), Very Large (XL), and World (W). The second set of instances, summarized in Table 4, is used for performance evaluation. It consists of five different World instances.
(a) Sites
(b) Graph sizes
(c) Flows
Figure 17: Description of scalability test instances
Table 4: Description of performance test instances
$\mathscr{G}$ nodes $\mathscr{G}$ arcs $\mathcal{G}$ nodes $\mathcal{G}$ arcs Bundles Orders Commodities
Min 19 689 160 466 63 518 1 192 022 7360 118 611 625 166
Max 19 769 161 024 63 778 1 198 366 7473 137 506 724 454
Benchmarks
We compare the performance and scalability of our ILS algorithm against the following benchmark heuristics:
Renault-based (R)
The current transport operations are projected into our modeling framework to produce an admissible solution.
Shortest (S)
Each bundle is assigned the shortest path in terms of distance.
Average (A)
All orders of a bundle are aggregated into a single order delivered at time step 1. The delivery path is then computed iteratively using the full giant container relaxation.
Constructive (C)
The constructive heuristic described in Section 4.1.
Lower Bound Rounding (LBR)
The rounding heuristic described in Section 5.
The Renault-based solution provides the first natural benchmark. As it reflects the industrial reality and not the planning strategy used, it acts as a proxy for the current planning strategy. The Shortest and Average heuristics represent conventional strategies commonly applied in practice. For completeness, we also include the constructive heuristic and the lower bound rounding heuristic. All benchmarks were implemented with identical code optimizations whenever applicable.
Because the model introduced in this paper is new to the literature, no directly comparable algorithm is available. On large-scale instances, the lower bound rounding heuristic serves as a natural mathematical programming baseline, consistent with rounding approaches used in flow problems (Jarrah et al., 2009; Lamothe et al., 2021; Lienkamp and Schiffer, 2024).
Implementation
We developed all algorithms in Julia (Bezanson et al., 2017) and released them as a public package, STPP-ILS.jl. To handle large-scale instances efficiently, we relied on several external libraries: Graphs.jl (Fairbanks et al., 2021) for custom graph structures and Dijkstra’s algorithm, OhMyThreads.jl (Bauer et al., 2025) for parallelization, and JuMP (Lubin et al., 2023) with Gurobi for MILP solving. All experiments were conducted on a machine equipped with an Intel Core i9 processor (2.20 GHz) and 64 GB RAM.
6.2 ILS hyper-parameters and time profiles
Several hyperparameters guide the execution of the ILS algorithm, as described in Section 4. The most important are the time limits for the local search and the perturbation phase. We calibrate perturbations further using two parameters: the maximum number of MILP variables and the minimum number of candidate paths. Figure 18 illustrates the influence of these parameters on both components and the resulting profile of th ILS obtained. Figure 18(a) shows that the local search start to converge after approximately 45 minutes. We therefore set the time limit for each local search in the ILS to 45 minutes. To balance the scale of perturbations with the degradation they introduce, we allocate two minutes for each perturbation. We restrict flow-based perturbations to 2,000,000 variables and require path-based perturbations to affect at least 30% of the bundles. We repeat perturbations until they modify 15% of the paths or increase the solution cost by more than 2%. Figure 18(c) reports the time profile of the ILS with these parameter settings on the World instance. As can be seen, the ILS converges after six hours of computation, corresponding to roughly six full iterations, and continues to improve only marginally thereafter. Accordingly, we adopt a six-hour time limit for all subsequent experiments in the computational study.
(a) Local search profile
(b) Perturbation impact
(c) ILS profile
Figure 18: Influence of parameters on local search and perturbation and the resulting ILS
6.3 Performance analysis
We assess how well the proposed ILS algorithm improves over Renault’s current planning solution and standard heuristic benchmarks, focusing on both scalability and solution quality. To do so, Figure 19 compare the performances of the introduced benchmarks against our ILS algorithm on the first set of instances. Additionally, Figure 20 compare the performances of the introduced benchmarks against our ILS algorithm on the second set of instances. It shows the time taken for each heuristic and the gap to the best lower bound available.
(a) Time
(b) Gap
(c) Gap (zoom)
Figure 19: Results on scalability test instances
Among all methods, the Iterated Local Search (ILS) achieves the highest solution quality, significantly outperforming all heuristic approaches, albeit at the cost of increased computational time. This additional runtime becomes increasingly justified as instance size grows: while ILS matches the performance of the greedy heuristic on small instances, despite being approximately a hundred times slower, it demonstrates clear performance advantages on larger instances.
Result 1.
On the World instances, the Iterated Local Search performs best, with an average gap to the lower bound of 7.9%. Further, it reveals a 23.2% potential for Renault’s operational plans cost reduction.
Result 2.
On the World instances, the Constructive heuristic performs second best, with an average gap to the lower bound of 9.9%. It is 30% faster than the lower bound rounding heuristic due to better parallelization.
(a) Time
(b) Gap
(c) Gap (zoom)
Figure 20: Results on performance test instances
Depending on the context, one will rather choose the ILS or the constructive heuristic. Strategic planning operations of large carriers such as Renault will choose the ILS, as every fraction of percentage point represents significant cost savings ($\sim$ 4 M€ per percent in our case). If faster evaluations are needed, like in tender calls, the constructive heuristic is recommended. Beyond ILS and the constructive heuristic, it is instructive to compare against simpler industry-style baselines, namely the Shortest and Average heuristics. These conventional heuristics deliver solutions quickly and scale well to large instances, but their neglect of consolidation limits solution quality. They therefore provide useful baselines, yet consistently fall short compared to tailored approaches such as ILS or the constructive heuristic.
Result 3.
Consolidation is necessary as illustrated by the poor performance of the Shortest heuristic. Average performs substantially better but is still 8.2% worse than the ILS.
6.4 Managerial and Practical Analyses
In addition to validating the performance of our algorithm, we conduct further analyses to derive insights that are directly relevant for practitioners. These analyses focus on three key aspects of transportation planning: the role of consolidation, transport regularity and outsourcing decisions. Each of these dimensions reflects a fundamental design choice in Renault’s inbound supply chain and influences both computational tractability and managerial decision-making.
Bin-packing consolidation.
Table 5 reports the three lower bounds on instance Medium, together with their associated solutions obtained by the rounding heuristic. As expected, the full giant container bound dominates the other bounds, followed by the mixed giant container and the linear relaxation. In practice, the mixed giant container bound emerges as the bound of choice: it is nearly as tight as the full giant container bound while requiring similar computation time as the linear relaxation. Importantly, the quality of the rounding heuristic remains poor, showing that stronger bounds do not automatically yield better feasible solutions. The gap between rounded solutions and the ILS solution is substantial: roughly 50% for the linear relaxation and 20% for the giant container bound. These results confirm that bin-packing is not only a computational bottleneck but also a central driver of solution quality.
Table 5: Comparison of lower bounds and associated rounding solutions on instance Medium
Bound Type Bound to bound gap Sol. to bound gap Sol. to sol. gap Time
Linear Relaxation -13.52% 64.50% 47.91% 1.0s
Mixed-Giant Cont. -0.73% 32.54% 19.17% 1.1s
Full-Giant Cont. 0.0% 35.31% 21.66% 3600s
Note : The full-giant container bound computation was stopped after 1 hour. It provides a worst solution than the mixed-giant because it tends to fill the giant-container approximation, leading to excess bin with a real bin packing.
Managerial insight: A faithful representation of consolidation is essential in industrial transportation planning. Approximating bin-packing by aggregated capacities, as often done in the literature, can lead to substantial distortions in solution quality. Our results suggest that even a one-dimensional bin-packing formulation, as used here, provides a meaningful and computationally tractable representation of consolidation effects.
Transport regularity.
Figure 21 shows the impact of time regularity on the World instance for both the constructive and the ILS heuristic.
(a) Instance size
(b) Constructive
(c) ILS
Figure 21: Impact of time regularity on the World instance
Lower levels of regularity cause a rapid increase in instance size, thereby escalating computational complexity. The constructive heuristic performance is heavily influenced by the level of regularity imposed. Reduced regularity degrades solution quality by up to 7%. In contrast, the ILS can exploit additional flexibility to achieve cost reductions of up to 1.05%. Interestingly, this increased flexibility has only marginal impact on the lower bound, which decreases by at most 0.25%, consistent with the near-linear nature of the relaxation.
Managerial insight: Regularity is highly valued by supply chain planners because it simplifies execution and reduces operational uncertainty. Our results confirm this preference: highly regular plans are not only easier to compute but also competitive in cost terms. While flexibility may yield improvements, the managerial trade-off favors regular transportation plans as a first step, especially in high-frequency inbound logistics.
Transport outsourcing.
Figure 22 compares the cost of a fully outsourced network to Renault’s current network across different volume configurations for the Medium and World instances. The fully outsourced case is obtained by replacing all non-outsourcing arcs with outsourcing costs estimated from the average rates observed in practice.
(a) Medium
(b) World
Figure 22: Impact of outsourcing on Medium and World instances
Our analysis reveals that pure outsourcing is suboptimal for high-volume contexts such as the automotive industry. Outsourcing remains attractive at low volumes, but as volumes increase, in-house planning becomes consistently superior. Larger networks reach the tipping point at which internal transportation planning becomes advantageous more quickly, because consolidation opportunities and economies of scale can be better exploited.
Managerial insight: The decision between outsourcing and internal planning depends critically on shipment volumes and network scale. For smaller networks or low-volume settings, outsourcing may be cost-effective. In contrast, for global, high-volume supply chains such as Renault’s, insourcing yields substantial savings and greater strategic control. Practitioners should therefore consider outsourcing only selectively, while maintaining in-house planning capabilities for the high-volume backbone of the network.
Summary.
In summary, these analyses highlight three central lessons for both researchers and practitioners. First, accurate modeling of consolidation is indispensable. Simplified representations of bin-packing may yield computational speed, but they introduce distortions that can misguide planning decisions. Even a one-dimensional formulation, as adopted here, provides the fidelity needed to capture Renault’s inbound consolidation challenges at scale. Second, regularity emerges as a double-edged design lever: from a computational perspective, highly regular plans reduce problem size and enable faster optimization; from a managerial perspective, they align with planners’ preference for predictable, stable operations. While relaxing regularity can unlock marginal cost improvements, the resulting complexity and planning uncertainty suggest that high-regularity plans constitute a robust and practical baseline. Third, outsourcing proves to be context-dependent. Although outsourcing is attractive in small networks or low-volume flows, large-scale inbound logistics—as faced by Renault—benefit considerably from in-house planning and execution. Internal planning not only captures economies of scale and consolidation benefits but also offers greater strategic control.
Taken together, these insights demonstrate that methodological advances such as ILS are not purely of academic interest: they directly inform the design of shipper transportation networks in practice. By clarifying when and why consolidation, regularity, and outsourcing matter, our results provide actionable guidance for managers seeking to balance cost efficiency, operational stability, and strategic flexibility in global supply chains.
7 Conclusion
This paper studied the shipper-side design of large-scale inbound transportation networks, motivated by Renault’s global supply chain. We introduced the Shipper Transportation Design Problem, which captures consolidation, routing, and regularity aspects, and developed a novel Iterated Local Search (ILS) metaheuristic to address it. The algorithm combines large-neighborhood search with MILP-based perturbations and leverages bundle-specific decompositions and giant container relaxations to generate tractable lower bounds and effective rounding benchmarks. Our computational study on real industrial data demonstrates that the proposed approach significantly improves upon both Renault’s current solutions and established benchmark heuristics. On the largest world-scale instances with more than 700,000 commodities and 1,200,000 arcs, the ILS achieves an average gap of 7.9% to the best available lower bound and delivers solutions showing a potential of 23.2% cost reduction compared to the Renault-based benchmark. To the best of our knowledge, this is the first work to successfully solve a shipper-side transportation design problem at such scale. Beyond methodological advances, our analysis provides clear managerial guidance. We show that accurate bin-packing models are indispensable to capture consolidation effects; that highly regular transportation plans not only reduce computational complexity but also align with planners’ preference for operational stability; and that outsourcing is only beneficial in low-volume contexts, while global high-volume supply chains benefit from in-house planning. These insights strengthen the case for advanced optimization methods in strategic shipper-side planning. In conclusion, this paper contributes a novel scalable algorithmic framework, the first computational evidence on global industrial-scale instances, and managerial lessons that directly support strategic decision-making in inbound logistics.
Future research could extend this work in several directions. Promising avenues include multi-dimensional consolidation models, dynamic outsourcing strategies that adapt to uncertain demand, and hybrid approaches that integrate predictive models with optimization for improved scalability. Such extensions would further enhance the applicability of decision-support tools for large industrial supply chains.
Appendix A Discussions and Extensions
As discussed in Section 3, the Shipper Transportation Planning model accommodates a wide range of extensions. Those are described in the following paragraphs.
Multimodal Transport
Few modes of transportation are used in Renault’s problem : trucks for inland transportation and boats for oversea transportation. Other modes could be used, such as trains, barge or planes, to carry parts. To take this into account, one could allow multiple arcs between two nodes, allowing for each arc to model one mode of transportation with different characteristics. Another option is to add add dummy nodes and arcs into the network. Figure 23 illustrates those two options.
P1P2Mode 1Mode 2Mode 3P1P2Mode 1Mode 2Mode 3
Figure 23: Example of multi-modal representation
Vehicle tours
Explicitly modeling vehicle tours, or milk-runs in the supply jargon, between suppliers and plants is not included in the scope of this work. They can however be approximated without the need to change our model by adding dummy platforms between proximate suppliers or plants. Collecting or delivery arcs from this platform would be free while inter-platform arcs would cost the full tour. Figure 24 illustrates this option.
S1S2S3US1S2S3PU
Figure 24: Example of vehicle tour approximation
Inventory
One can also notice that no mention of inventory is being made throughout the model. This stems in our case from the coarse time step taken, making all inventory happens inside time steps, which make them invisible. Finer temporal resolution enables explicit inventory modeling by adding intra-node arcs between successive time steps, called inventory arcs. An adequate inventory cost would be applied on those arcs. Figure 25 illustrates this.
P1P1P1T = 1T = 2T = 3Part SupplierS1S2S3
Figure 25: Example of inventory (left) and part sourcing (right) representation
Part Sourcing
As previously mentioned, in our context, the same part can be provided by several suppliers to several units but they are not considered substitutes and are treated mathematically as different commodities. Integrating part sourcing into this model can be done by adding a dummy node for each part, like a super-supplier $S^{\mathrm{part}}$, connected to all eligible suppliers of this part. Figure 25 illustrates this.
The flow constraints on $\mathcal{G}$ would then need to be explicitly stated , as followed :
$\sum_{a\in\delta^{+}(\nu)}f_{a}^{m}-\sum_{a\in\delta^{-}(\nu)}f_{a}^{m}=e_{\nu}^{m}\;\;\forall m\in M\text{ , }\nu\in\mathcal{V}\cup S^{m}$
$f_{\pi_{\mathcal{G}}(\alpha,o)}^{m}\leq C^{m}x_{\alpha}^{b}\quad\forall b\in B\text{ , }o\in O_{b}\text{ , }m\in M_{o}$
Return logistics
By considering packaging used in plants as commodities to be due to suppliers, adding a dummy supplier for each plant and a dummy plant for each supplier, this model can tackle simultaneously forward and return logistics of parts.
Appendix B Proof of Proposition 2
We consider the linear relaxation of problem (9). For $a\in\mathcal{A}^{\mathrm{con}}$ and $k\in K_{a}$, as $\tau_{a}^{k}\in\mathds{R}_{+}$, we have :
$L_{a}\tau_{a}^{k}=\sum_{m\in M}y_{ak}^{m}\ell_{m}$
For $a\in\mathcal{A}^{\mathrm{con}}$ and $m\in M$, bin packing constraints imposes :
$\sum_{k\in K_{a}}y_{ak}^{m}=f_{a}^{m}$
We have therefore for all $a\in\mathcal{A}^{\mathrm{con}}$ :
$\sum_{k\in K_{a}}\tau_{a}^{k}=\sum_{m\in M}f_{a}^{m}\frac{\ell_{m}}{L_{a}}$
In this relaxation, the number of transport units used by commodities on each arc equals the total commodity volume on the arc divided by the unit capacity. These equations transforms packing constraints into “liquid container” volume cost, accounted for in the linearly relaxed network cost function in problem (16).
$\sum_{a\in\mathcal{A}}\sum_{m\in M}(c_{am}^{\mathrm{com}}+\frac{\ell_{m}}{L_{a}}c_{a}^{\mathrm{con}})f_{a}^{m}+\sum_{p\in\mathcal{P}}c_{p}^{\mathrm{over}}z_{p}$
Without platform overloading constraints (7), this linear relaxation of bin-packing constraints allows decoupling between commodities, which is equivalent in our case to a decoupling between bundles. Our shortest path constraint is written as a flow constraint on the graph $\mathscr{G}$. Because the polytope of flows is perfect, the value we obtain is the linear relaxation bound.  $\square$